OpenStudy (anonymous):

Collect like terms: x^2+2xy+y^2=x^2-y^2

OpenStudy (anonymous):

we factor the right side i think (x+y)^2 = (x+Y)(x-y)

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

go on chat

OpenStudy (anonymous):

2xy + 2y^2 =0

OpenStudy (anonymous):

subtract both sides x^2